OpenStudy (anonymous):

X number of men finish a work in 60 days. If there were 8 more men, the work would be done in 10 days less. Find X.

OpenStudy (anonymous):

Why doesn't this work? --- [(X+8)/60]= 1/50 ?

OpenStudy (anonymous):

Right now, you are saying ...uhm... like x+8 men can finish the work in 60 days, or something like that. x men finished the work in 60 days. If there were x+8 men, the work can be done in 10 days less, so 50 days. The correct proportion is (x/60) = ((x+8)/50). Find x.

OpenStudy (anonymous):

ah! thank you!

OpenStudy (anonymous):

The previous reply is incorrect. Let w = the amount of work x = the number of men doing the work Then 60x = w and 50(x+8) = w. Setting both equations equal to each other: 60x = 50(x+8) or 60x = 50x + 400 so 10x = 400 or x = 40

OpenStudy (anonymous):

Aw, I'm so sorry :(